One number is 11 more than another. Find the two numbers if three times the larger number exceeds four times the smaller number by 4.
Found 2 solutions by mananth, nyc_function:
Answer by mananth(16949)

One number is 11 more than another. Find the two numbers if three times the larger number exceeds four times the smaller number by 4.
..
let larger number be x
the smaller number be y
..
x-y = 11...............1
3x = 4y +4
3x-4y=4................2
..
multiply equation 1 by -3 and add it to equation 2
-3x+3y=-33
3x-4y=4
-y=-29
y= 29
..
x-y = 11
x-29=11
x= 29+11
x= 40
The numbers are 29,40

Answer by nyc_function(2741)

Let x = smaller number
Let x + 11 = bigger number
3(x + 11) = 4x + 4
3x + 33 = 4x + 4
3x - 4x = 4 - 33
-x = -29
x = -29/-1
x = 29...This is the smaller number.
29 + 11 = 40...This is the bigger number.